# Disentangling Memory and Reasoning Ability in Large Language Models

### v2

@@ Abstract
Large Language Models (LLMs) have demonstrated strong performance in handling complex tasks requiring both extensive knowledge and reasoning abilities. However, the existing LLM inference pipeline operates as an opaque process without explicit separation between knowledge retrieval and reasoning steps, making the model’s decision-making process unclear and disorganized. This ambiguity can lead to issues such as hallucinations and knowledge forgetting, which significantly impact the reliability of LLMs in high-stakes domains. In this paper, we propose a new inference paradigm that decomposes the complex inference process into two distinct and clear actions: (1) memory recall: which retrieves relevant knowledge, and (2) reasoning: which performs logical steps based on the recalled knowledge. To facilitate this decomposition, we introduce two special tokens and , guiding the model to distinguish between steps that require knowledge retrieval and those that involve reasoning. Our experiment results show that this decomposition not only improves model performance but also enhances the interpretability of the inference process, enabling users to identify sources of error and refine model responses effectively. The code is available at https://github.com/MingyuJ666/Disentangling-Memory-and-Reasoning.
@@ 1 Introduction
Recent advancements in Large Language Models (LLMs) have showcased their impressive inference capabilities in handling complex natural language tasks that require both extensive knowledge and sophisticated reasoning abilities (OpenAI, 2024; Touvron et al., 2023; Wei et al., 2022a). LLMs have demonstrated the ability to memorize vast amounts of knowledge, and techniques like Chain-of-Thought (CoT) (Wei et al., 2022b), Tree of thoughts (ToT) (Yao et al., 2024) have been developed to further enhance their inference abilities by decomposing complex problems into several simpler, single-step processes. These methods enable LLMs to tackle multi-step inference tasks more effectively by organizing the thought process into discrete, focused actions (Feng et al., 2024; Jin et al., 2024b; Wei et al., 2022b).
@@
However, despite these advancements, existing inference frameworks often operate as an opaque process without explicit separation between knowledge retrieval and reasoning steps. This makes it unclear what specific knowledge the model utilizes and how it performs reasoning, leaving the decision-making process ambiguous. For complex, knowledge-intensive tasks, such as multi-hop inference, LLMs often struggle to effectively leverage their memory for inference (Yang et al., 2023; Jin et al., 2024b; Wang et al., 2024b; Cheng et al., 2024; Liu et al., 2024). Such tasks typically require the ability to recall relevant knowledge for each reasoning step (or "hop") and then perform inference over that recalled memory (Wang et al., 2024c). The lack of structure in the output and effective memory utilization can lead to issues such as hallucinations, where LLMs generate plausible but incorrect information (Xu et al., 2024; Li et al., 2024a), and "forgetting," where relevant information is lost across reasoning steps (Jin et al., 2024b; Chen & Shu, 2023), disrupting the logical flow. These issues compromise the LLM’s accuracy and reliability, posing serious risks in high-stakes applications like healthcare and finance Pham & Vo (2024).
@@
Existing works on enhancing inference in LLMs to address their challenges can be categorized into two approaches: Memory-Based Approaches: These methods focus on improving the recall and utilization of world knowledge stored in the model, such as leveraging Retrieval-Augmented Generation (RAG) (Cai et al., 2019; Chen et al., 2024b). The emphasis is on enabling models to access and use their outside knowledge more effectively. Reasoning-Based Approaches: These techniques aim to improve the reasoning capabilities of models by Chain-of-Thought (CoT) reasoning (Yang et al., 2023; Gao et al., 2024; Yu et al., 2024) or introducing structured guidance in training such as planning tokens (Wang et al., 2023b, 2024b) to organize reasoning into discrete, interpretable steps. These methods enhance the ability of LLMs to handle complex reasoning tasks by embedding structural reasoning mechanisms into their parameters. Despite advancements in both categories, LLMs still struggle with tasks requiring an intricate interplay of memory recall and logical reasoning (Wang et al., 2024c).
@@
In this work, we introduce a new LLM inference paradigm that decomposes the complex inference process into two distinct actions: memory and reasoning. To implement this paradigm, we first construct a new training dataset by leveraging a powerful LLM such as GPT-4o to generate itemized actions for various question-answering tasks, with each action labeled as either reason or memory. Each action is preceded by a special token, either or , which serves as a control signal during training. The second step involves training a customized LLM using this labeled dataset. The inclusion of these two learnable control tokens guides the model to distinguish between recalling relevant knowledge and performing reasoning steps. With these control tokens, we encourage the trained model to first use memory to recall relevant knowledge and then apply reasoning based on that memory to answer the question. This approach not only constrains the model’s behavior but also enhances the interpretability of the inference process.
@@
Our experimental results demonstrate that the proposed decomposition improves performance and enhances the interpretability of the model’s inference process. Specifically, our method achieves accuracy of 78.6% and 78.0% on the StrategyQA dataset (Geva et al., 2021) using Qwen2.5-7B (Yang et al., 2024a) and LLaMA-3.1-8B (Touvron et al., 2023), respectively. These results represent improvements of 1.2% and 1.3% over the planning-token fine-tuned baseline while remaining only 2.2% below GPT-4o’s performance. Remarkably, on the TruthfulQA dataset (Lin et al., 2022), LLaMA-3.1-8B enhanced by our algorithm outperforms GPT-4o with Chain of Thought prompting (85.4%), achieving 86.6% accuracy. On average across three benchmark datasets, our method narrows the performance gap with the top-performing closed-source model, GPT-4o (using CoT prompting), to just 1.9%. Furthermore, our algorithm achieves a disentanglement accuracy of 83.3%, providing insights into the model’s inference. Analyzing the errors made by LLaMA-3.1-8B reveals that most issues stem from reasoning, rather than deficiencies in the knowledge itself. This distinction sheds light on the primary sources of errors in the model’s outputs and enables targeted improvements. Notably, our evaluation approach incorporates a detailed error ratio analysis that distinguishes between memory and reasoning steps, offering a more nuanced understanding of model limitations compared to traditional accuracy metrics. By identifying specific areas for refinement, this analysis supports more effective and informed enhancements to model performance.
@@
Improve Benchmark Performance: Our model achieves competitive results, surpassing GPT-4o on TruthfulQA and closely matching GPT4-o on StrategyQA and CommonsenseQA, demonstrating the benefits of our approach.
Enhance Interpretability and Control: Our framework enables transparent reasoning with labeled steps for memory and reasoning, allowing precise error analysis and control in high-stakes applications.
@@ 2 Method
The workflow of our algorithm is illustrated in Figure 2. Given a question, we first adopt an LLM generation framework, consisting of a knowledge LLM to extract relevant knowledge, followed by a reasoning LLM to derive logical steps. The data we generated for the question are annotated in steps of and . Finally, the questions with annotated steps and answers are then combined to train the model, where the annotations serve as a special trainable token.
2.1 Data Generation
@@
In this section, we introduce an LLM-based framework to generate memory (knowledge in LLM) and reasoning steps, consisting of an inference LLM and a knowledge LLM, as illustrated in Figure 2. First, we use an inference LLM to generate Chain of Thought (CoT) (Wei et al., 2022b) inference steps, prompting it to mark steps that require factual knowledge as and those requiring reasoning as . To improve the quality of the memory steps, we further instruct the inference LLM to rephrase knowledge marked as into questions, emphasizing its factual nature. For the example question in Figure 1, The inference LLM first retrieves relevant knowledge about MMA and Roman Colosseum games, analyzes their relationship and synthesizes this information to form a coherent judgment. By methodically aligning each step with its purpose, the LLM ensures that the conclusion—MMA is not "totally original" from the Colosseum games—reflects well-supported reasoning. Next, a knowledge LLM answers the additional factual questions generated by the inference LLM, such as What is mixed martial arts? and What are the Roman Colosseum games?. The answers to these questions are then substituted into the inference steps. This approach not only effectively decouples reasoning and knowledge to ensure accuracy but also maintains high data quality. This data generation approach is designed to fine-tune LLMs so that LLMs can effectively disentangle knowledge and reasoning during inference.
2.2 Training
In this method, we train a LLM by incorporating intervened reasoning and memory processes as Figure 2, guided by two distinct tokens: , which represents reasoning with knowledge, and , which signifies retrieved knowledge. These tokens are designed to prompt the model to activate specific reason and knowledge abilities, strengthening its reasoning capabilities and ultimately enhancing both interpretability and performance in complex inference tasks.
@@ 2.2 Training
In training time, each training instance comprises the following components: (1) the question tokens where is the question token length, (2) the step-by-step thinking process consists of intertwined memory and reasoning elements, denoted as and , where each is initiated by a special token followed by a sequence of tokens that represent retrieved factual or stored knowledge: , and is initiated by a special token followed by a sequence of tokens that represent the reasoning process: , and (3) the target answer generated after the completion of the memory retrieval and reasoning processes. The model is trained in a standard autoregressive manner using LoRA fine-tuning and the and are trainable out of vocabulary tokens.
@@
By structuring the input in this manner, the model learns to process and distinguish between retrieved knowledge and the reasoning steps required to generate the final answer. The inclusion of the and tokens facilitates the disentanglement of memory retrieval and reasoning processes, thereby enhancing the model’s ability to produce coherent and accurate responses.
@@ 3.1 Experiment Setup
In our experiments, we use LLaMA-2-7B-chat-hf (Touvron et al., 2023), LLaMA-3.1-8B-Instruct (Dubey et al., 2024), and Qwen2.5-7B-Instruct (Yang et al., 2024a) as target models for training. GPT-4o serves as the inference and knowledge LLM to generate training data, while GPT-4o-mini is employed as the evaluator.
@@
Our experiments are conducted on three datasets. StrategyQA (Geva et al., 2021) is a question-answering benchmark comprising 2,780 examples that require inference, with each example containing a strategy question, a Chain-of-Thought (CoT) path, and supporting evidence. CommonsenseQA (Talmor et al., 2019) is a question-answering dataset with 12,102 questions, each necessitating commonsense knowledge to identify the correct answer among four distractors. TruthfulQA (Lin et al., 2022) serves as a benchmark for evaluating the truthfulness of language model responses, encompassing 817 questions across 38 categories. For TruthfulQA, we use the mc1_targets, which contains single-choice questions with 4-5 answer choices per question. We transformed this dataset into a single-choice format to facilitate training and testing with answer options labeled as A-E. In the original dataset, the correct answer was always the first option. We shuffled the labels to prevent the model from learning shortcuts. StrategyQA and CommonsenseQA datasets have predefined train and test splits for mc1_targets in TruthfulQA, which we divided into an 8:2 ratio for the training and testing sets.
@@
Five main methods are being compared: Zero-shot, CoT (Chain-of-Thought), LoRA, and Planning-token. In our experiments, we adopt zero shot (Just input the question) and Chain of Thought (CoT) prompting as our vanilla baseline for reasoning. For the fine-tuned baseline, we use int8_training to save memory and we choose LoRA fine-tuning (Hu et al., 2021) and Planning Tokens (LoRA+Prompt Tuning) (Wang et al., 2023b) to train and test on these three datasets to facilitate a comparative evaluation with our proposed approach.
@@
We use accuracy (acc) to measure the model’s performance on multiple-choice questions.
@@ 3.2 Main Results
Our algorithm consistently achieves superior performance across multiple benchmarks compared to existing approaches. On StrategyQA, LLaMA-3.1-8B fine-tuned by our method achieves 78.0%, outperforming CoT (69.4%) and Planning-token (76.7%), while LLaMA-2-7B fine-tuned with our method achieves 70.6%, compared to LoRA (61.2%) and Planning-token (63.5%). Similarly, in CommonsenseQA, "Ours" with LLaMA-3.1-8B scores 82.3%, surpassing CoT (70.6%) and Planning-token (76.9%), while LLaMA-2-7B with our method achieves 71.1%, outperforming LoRA (65.4%). A significant breakthrough is observed on the TruthfulQA dataset, where LLaMA-3.1-8B enhanced by our method achieves 86.6%, outperforming GPT-4o by 1.8% in zero-shot setting and 1.2% in CoT setting, and also surpassing Planning-token (82.5%) and LoRA (79.8%) in fine-tuned settings. Notably, on StrategyQA, Qwen2.5-7B fine-tuned with our method achieves 78.6%, exceeding Planning-token performance (77.4%) and LoRA (69.1%). On the TruthfulQA dataset, Qwen2.5-7B struggles with CoT settings, achieving only 56.7%, which underscores the instability of instruction-tuned models like Qwen2.5-7B on certain datasets. These results emphasize the robustness and adaptability of our approach in handling inference-intensive tasks across diverse datasets and models. However, adding a CoT inference step can sometimes decrease performance, as also reported by Sprague et al. (2024). The unstable training of instruction-tuned models like Qwen2.5-7B further highlights the limitations of existing methods compared to our approach.
@@
3.3 Ablation Study
3.3.1 The contribution of Memory and Reason token?
The Ablation Experiment presents an ablation study comparing the performance of two versions of the LLaMA model ( LLaMA-2-7B and LLaMA-3.1-8B) across three benchmarks: StrategyQA, CommonsenseQA, and TruthfulQA. The study examines the impact of using specific tokens ("Memory and Reason" vs. "Random") on the model’s performance. During training, we shuffled the allocation of and tokens and then observed the effects on training and testing performance. As expected, the overall performance declined, but the rate of decline varied, demonstrating the superiority of our approach in disentangling reason and memory.
3.3.2 The impact of the number of special tokens
In our training setup, in addition to the two token types, and , we include a parameter representing the number of special tokens preceding each sentence. A sentence might include three tokens or four like the example above, with the same variation applied to memory tokens. Our experiments indicate that model performance improves with around four to six special tokens, likely because more tokens can lead to better performance for the model during training (Levy et al., 2024). As shown in table 4 and table 4, we selected two models ( LLaMA-2-7B, LLaMA-3.1-8B) to illustrate their performance across different numbers of special tokens.

### v3

@@ Abstract
Large Language Models (LLMs) have demonstrated strong performance in handling complex tasks that require both extensive knowledge and reasoning abilities. However, the existing LLM inference pipeline operates as an opaque process without explicit separation between knowledge retrieval and reasoning steps, making the model’s decision-making process unclear and disorganized. Recent research has shown that this ambiguity will lead to issues such as knowledge forgetting, which significantly impact the reliability of LLMs. In this paper, we propose a novel language model inference paradigm that decomposes the complex inference process into two distinct and clear actions: (1) memory recall: which retrieves relevant knowledge in LLM, and (2) reasoning: which performs reasoning steps based on the recalled knowledge. To facilitate this decomposition, we introduce two special tokens and , guiding the model to distinguish between steps that require knowledge retrieval and those that involve reasoning. Our experiment results show that this decomposition not only improves LLMs’ performance among utility benchmarks but also enhances interpretability during the inference process, enabling users to identify sources of error and refine model responses effectively. The code is available at: https://github.com/MingyuJ666/Disentangling-Memory-and-Reasoning.
Disentangling Memory and Reasoning Ability in Large Language Models
Mingyu Jin1 Weidi Luo2 Sitao Cheng3 Xinyi Wang3 Wenyue Hua3 Ruixiang Tang1 William Wang3 Yongfeng Zhang1††thanks: Corresponding Email: mingyu.jin@rutgers.edu yongfeng.zhang@rutgers.edu. 1Rutgers University 2The Ohio State University 3University of California, Santa Barbara
@@ 1 Introduction
Recent advancements in Large Language Models (LLMs) have showcased their impressive inference capabilities in handling complex natural language tasks that require both extensive knowledge and sophisticated reasoning abilities OpenAI (2024); Touvron et al. (2023); Wei et al. (2022a). LLMs have demonstrated the ability to memorize vast amounts of knowledge, and techniques like Chain-of-Thought (CoT) Wei et al. (2022b), Tree of thoughts (ToT) Yao et al. (2024) have been developed to further enhance their inference abilities by decomposing complex problems into several simpler, single-step processes. These methods enable LLMs to tackle multi-step inference tasks more effectively by organizing the thought process into discrete, focused actions Feng et al. (2024); Jin et al. (2024); Wei et al. (2022b); Sun et al. (2025).
@@
Despite these advancements, existing inference frameworks often operate as an opaque process without explicitly separating knowledge retrieval and reasoning steps. This makes it unclear what specific knowledge the model utilizes and how it performs reasoning, leaving the decision-making process ambiguous. For complex, knowledge-intensive tasks, LLMs often struggle to effectively leverage their memory for inference Yang et al. (2023); Jin et al. (2024); Cheng et al. (2024); Liu et al. (2024). Such tasks typically require the ability to recall relevant knowledge for each reasoning step and then perform inference over that recalled memory Wang et al. (2024c). The lack of structure in the output and the inefficient memory utilization can result in issues such as knowledge forgetting, where relevant information is lost across reasoning steps Chen and Shu (2023), which disrupts the logical flow, as well as hallucinations, where LLMs generate plausible yet incorrect information Xu et al. (2024); Li et al. (2024a). These issues compromise the LLM’s accuracy and reliability, posing serious risks in high-stakes applications like healthcare and finance Pham and Vo (2024).
@@
Existing efforts to enhance inference in LLMs and address their challenges can be broadly classified into two main approaches: Memory-based Approaches: These methods focus on improving the recall and utilization of world knowledge that may not be stored in the model, such as leveraging Retrieval-Augmented Generation (RAG) Cai et al. (2019); Chen et al. (2024b). The emphasis is on enabling models to access and use their outside knowledge more effectively. Reasoning-based Approaches: These techniques aim to improve the reasoning capabilities of models by Chain-of-Thought (CoT) reasoning Yang et al. (2023); Gao et al. (2024); Yu et al. (2024) or introducing structured guidance in training such as planning tokens Wang et al. (2024d, b) to organize reasoning into discrete, interpretable steps. These methods enhance the ability of LLMs to handle complex reasoning tasks by embedding structural reasoning mechanisms into their parameters. Despite advancements in both categories, LLMs still struggle with tasks that require an intricate interplay of memory recall and logical reasoning Wang et al. (2024c).
@@
In this work, we propose a novel LLM inference paradigm that divides the complex inference process into two distinct components: memory and reasoning. Specifically, we generate itemized action responses for various question-answering datasets, categorizing each action as either memory or reasoning. Each action is then preceded by a special token, either or , which acts as a control signal during training. The second step involves training an LLM using these modified outputs. By incorporating these learnable control tokens, the model is explicitly guided to distinguish between recalling relevant knowledge and performing reasoning steps. This structured guidance encourages the model to first use memory to retrieve the relevant information and then apply reasoning based on that memory to solve the task. Our approach not only introduces a new form of structured response generation but also establishes a novel framework for guiding LLMs to "think" systematically. This structured decomposition improves both the model’s performance and the interpretability of its inference process.
@@
Our experimental results demonstrate that the proposed decomposition improves performance and enhances the interpretability of the model’s inference process. Specifically, our method achieves accuracy of 78.6% and 78.0% on the StrategyQA dataset Geva et al. (2021) using Qwen2.5-7B Yang et al. (2024a) and LLaMA-3.1-8B Touvron et al. (2023), respectively. These results represent improvements of 1.2% and 1.3% over the planning-token fine-tuned baseline while remaining only 2.2% below GPT-4o’s performance. Remarkably, on the TruthfulQA dataset Lin et al. (2022), LLaMA-3.1-8B enhanced by our algorithm outperforms GPT-4o with Chain of Thought prompting (85.4%), achieving 86.6% accuracy. On average across three benchmark datasets, our method narrows the performance gap with the top-performing closed-source model, GPT-4o (using CoT prompting), to just 1.9%. Furthermore, by analyzing the errors made by LLaMA-3.1-8B, we reveal that most issues stem from reasoning rather than deficiencies in the knowledge itself. This distinction sheds light on the primary sources of errors in the model’s outputs and enables targeted improvements.
@@
Advancing Benchmark Performance: Our model achieves competitive results, surpassing GPT-4o on TruthfulQA and closely matching GPT4-o on StrategyQA and CommonsenseQA, demonstrating the benefits of our approach.
Empowering Transparency and Control: Our framework enables transparent reasoning with labeled steps for memory and reasoning, allowing precise error analysis and model refinement.
@@ 2 Method
The workflow of our method can be divided into two stages: Data generation by decoupling memory and reasoning steps and training LLM with memory and reasoning tokens on generated data.
2.1 Data Generation with Decoupled Memory and Reasoning
@@
We introduce an LLM-based framework for response generation to generate memory (knowledge in LLM) and reasoning steps, consisting of an inference LLM and a knowledge LLM, as illustrated in Figure 1. First, we use an inference LLM to generate Chain of Thought (CoT) (Wei et al., 2022b) inference steps, prompting it to mark steps that require factual knowledge as and those requiring reasoning as . To improve the quality of the memory steps, we further instruct the inference LLM to rephrase knowledge marked as into questions, emphasizing its factual nature. For the example question in Figure 1, The inference LLM first retrieves relevant knowledge about MMA and Roman Colosseum games, analyzes their relationship and synthesizes this information to form a coherent judgment. By methodically aligning each step with its purpose, the LLM ensures that the conclusion—MMA is not "totally original" from the Colosseum games—reflects well-supported reasoning. Next, a knowledge LLM answers the questions about factual knowledge generated by inference LLM, such as What are the origins and characteristics of mixed martial arts? and What were the Roman Colosseum games?. The answers to these questions are then substituted into the CoT inference steps. This approach effectively decouples reasoning from knowledge, ensuring accuracy while maintaining high data quality. It enables the fine-tuning of LLMs by disentangling knowledge and reasoning during inference. We leverage this LLM-based framework between memory and reasoning steps to generate interpretable data, which can be used for the training stage.
2.2 LLM Training with Memory and Reasoning Tokens
@@
At this stage, we train an LLM by incorporating intervened reasoning and memory processes as Figure 1, guided by two special tokens: , which represents reasoning with knowledge, and , which signifies retrieved factual knowledge. These special tokens are designed to prompt the model to activate the necessary knowledge for reasoning, strengthening its inference capabilities and ultimately enhancing both interpretability and performance in complex inference tasks. During training stage, each training instance comprises the following components: (1) the question tokens where is the question token length, (2) the step-by-step thinking process consists of intertwined memory and reasoning components, denoted as and , where each is initiated by a special token followed by a sequence of tokens that represent retrieved factual knowledge: , and is initiated by a special token followed by a sequence of tokens that represent the reasoning process: , and (3) the target answer generated after the completion of the memory retrieval and reasoning processes. The model is trained in a standard autoregressive manner using LoRA fine-tuning and the and are trainable out of vocabulary tokens.
@@
By structuring the input in this paradigm, the model learns to process and distinguish between retrieved knowledge and the reasoning steps required to generate the final answer. The inclusion of the and tokens facilitates the disentanglement of memory retrieval and reasoning processes, thereby enhancing the model’s ability to produce coherent and accurate responses.
@@ 3.1 Experiment Setup
In our experiments, we use LLaMA-2-7B-chat-hf Touvron et al. (2023), LLaMA-3.1-8B-Instruct Dubey et al. (2024), and Qwen2.5-7B-Instruct Yang et al. (2024a) as backbone models for training and test. GPT-4o serves as the inference and knowledge LLM to generate training data, while GPT-4o-mini is employed as the evaluator.
@@
Our experiments are carried out on three data sets: StrategyQA Geva et al. (2021) is a question-answer benchmark of 2,780 examples. Each example includes questions, supporting evidence, and answers. CommonsenseQA Talmor et al. (2019) contains 12,102 questions, each of which requires common sense knowledge to select the correct answer from four distractors. TruthfulQA Lin et al. (2022) evaluates the truthfulness of the responses to the language model, with 817 questions in 38 categories. We used the mc1_targets subset, which consists of single-choice questions with 4-5 answer choices. To prepare this dataset for training and testing, we labeled the answer options as A-E and shuffled the labels to avoid shortcuts in training. For StrategyQA and CommonsenseQA, we used their predefined training and testing set splits. For TruthfulQA, we split the data into an 8:2 ratio for training and testing.
@@
In our experiments, we adopt zero shot (just input the question) and CoT prompting as our vanilla baseline for inference. For the fine-tuned baseline, we choose LoRA fine-tuning Hu et al. (2021) and Planning Tokens (LoRA+Prompt Tuning) Wang et al. (2024d) to train and test on these three datasets to facilitate a comparative evaluation with our approach. In training, we use int8_training to save GPU memory and accelerate.
@@
We use accuracy (acc) to measure the model’s performance on all datasets.
@@ 3.2 Main Results
Five main methods are being compared: Zero-shot, CoT (Chain-of-Thought), LoRA, Planning-token, and Ours (mentioned in Section 3.1). The results in StrategyQA and CommonsenseQA benchmarks indicate that our algorithm consistently achieves higher scores across both benchmarks compared to other approaches, particularly in fine-tuned models. For instance, in StrategyQA, Our method enhanced LLaMA-3.1-8B achieved a score of 78.0%, outperforming CoT at 69.4% and Planning-token at 76.7%. Similarly, in CommonsenseQA, Our method enhanced LLaMA-3.1-8B scores 82.3%, compared to CoT’s 70.6% and Planning-token’s 76.9%, suggesting the effectiveness of our algorithm in improving LLMs’ performance.
For the TruthfulQA dataset, we achieved a significant breakthrough; the LLaMA-3.1-8B enhanced by our algorithm (86.6%) even outperforms GPT-4o in both zero-shot (84.8%) and CoT settings (85.4%), which is remarkable. GPT-4 sometimes gets misled by these options in this dataset, but our model effectively handles these challenges. Our model first considers relevant knowledge and then uses it in reasoning, which proves highly effective on this dataset(as the appendix E, we include an analysis of both correct and incorrect examples). However, Qwen2.5-7B performed poorly on this dataset, achieving only 81.0% in our algorithm, likely due to instruction tuning in Qwen2.5, resulting in average performance and unstable training. However, adding a CoT can decrease performance for some models in some datasets, which is also a phenomenon reported by Sprague et al. (2024).
@@
3.3 Ablation Study
In the ablation study, we comprehensively investigate the effects of the impact of special token 3.3.1 and the number of special tokens 3.3.2.
3.3.1 Impact of Memory and Reason Tokens
The Ablation Experiment presents an ablation study comparing the performance of two versions of the LLaMA model (LLaMA-2-7B and LLaMA-3.1-8B) across three benchmarks: StrategyQA, CommonsenseQA, and TruthfulQA. The study examines the impact of using specific tokens ("Memory and Reason") vs.Random tokens on the model’s performance. During training, we shuffled the allocation of and tokens and then observed the effects on training and testing performance. As expected, the overall performance declined (shown in Table 2), but the decline rate varied (from 2.1% to 6.6%), showing our approach’s superiority in disentangling reason and memory.
3.3.2 Impact of Special Token Count
In our training setup, we have two token types: and , and we will include a parameter representing the number of special tokens preceding each sentence. For example, A sentence might include three tokens or four, like the question in Appendix D.3. Our experiments indicate that model performance reaches a higher point with around four to six special tokens (as Table 4 and 4). This is likely because more tokens may lead to better performance for the LLM Levy et al. (2024), as proved by previous research. We selected two LLMs to illustrate their performance (ACC) across different numbers of special tokens.
Another important issue is knowledge distillation. We must ensure that the model’s improvement is not due to knowledge distillation from the GPT-4 framework. Using the same inference steps, we compared the results of standard training with 0 reason and memory tokens and found that adding these tokens significantly improves performance. This indirectly confirms that the model’s enhancement comes from algorithmic improvements rather than knowledge distillation.